OpenStudy (anonymous):

You have $60 and your sister has $90. You are planning to save $8 per week and your sister is planning to save $5 per week. How long will it take before you and your sister have the same amount of money?

OpenStudy (zpupster):

x= no of weeks 60+8x= amount+amount* weeks--you 90+5x= amount+amount* weeks--sister we want to know when they are equal 60+8x=90+5x can you solve for x

OpenStudy (anonymous):

it is going to take you X weeks 1. after x weeks you will be having$8 times xweeks plus what you already had,$60,8x +60 2.after x weeks sister will be having $5 times xweeks + what she already had,$90,5x +90 3 you will both be having the same amounts i.e , 8x + 60 =5x + 90 solve for x, to get the number of weeks!